# Domain-wall melting and entanglement in free-fermion chains with a band structure

### v1

@@ 2 Model and setup
We will be interested in the dynamics generated by the quadratic Hamiltonian (1). Assuming a Gaussian initial state |$\ket{\psi(0)}$, all the information about the time-evolved state is encoded in the two-point correlation matrix $C(t)$, with matrix elements given by
$C_{m,n}(t)=\bra{\psi(t)}c_{m}^{\dagger}c_{n}\ket{\psi(t)},\qquad\ket{\psi(t)}=\mathrm{e}^{-i\hat{H}t}\ket{\psi(0)}.$ (6)
@@ 3 Correlations and hydrodynamics
The result (12) has a very clear interpretation in terms of the quasiparticle picture. Indeed, the two terms simply correspond to contributions from the lower and upper band of quasiparticles $\alpha_{q}$ and $\beta_{q}$, respectively. Due to the symmetry of the dispersion in (4), these modes propagate ballistically with velocities ±$\pm v_{q}$, and contribute to the correlations if they reach the midpoint between the cells $m$ and $n$. Note also that (12) can be rewritten more compactly as
@@ 3.1 Density and current
We first evaluate the particle density and current defined in (8). Note that the result is going to depend only on $|\delta|$, and we assume $10<\delta<$ 1 to simplify notation. The maximal quasiparticle velocity can be obtained via the solution of $\frac{\mathrm{d}v_{q}}{\mathrm{d}q}=0$, which yields
@@ 4.1 CFT results
Having identified the inhomogeneous CFT, the next step is to apply the replica trick and twist-field machinery [48, 49] for the computation of the entropy of the domain $A=[-\infty,x]$, which has been carried out in [22]. The key is to map the inhomogeneous CFT (28), defined on a strip $[0,\pi]\times\mathbb{R}$ in terms of the isothermal coordinates (27), into a flat one living on the upper half plane. This can be achieved by the subsequent application of a Weyl-transformation $\mathrm{e}^{\sigma(z,\bar{z})}$, and the exponential map $g(z)=\mathrm{e}^{iz}$. The twist field inserted at $(x,y)$ in the original geometry is thus mapped into $g(z)$ on the flat upper half plane, which amounts to calculating the entropy of a segment of length $\mathrm{Im}\,g(z)$ located at the boundary of a half-infinite chain. Taking into account the Jacobians of the conformal transformations, one obtains an effective length scale and the universal contribution to the entropy is given by [22]
@@
However, this is not the end of the story, as we have to deal with a lattice problem, which introduces a space-dependent non-universal contribution to the entropy. Nevertheless, this can be found by applying a local density approximation and using the known results for the constant term in a homogeneous chain [50]. Indeed, the time-evolved state is locally described by a boosted Fermi sea $[k_{-},k_{+}]$ with boundaries $k_{\pm}=\frac{\pi}{2}\pm\arccos\nu$, such that the lattice contribution is given by
@@ 4.2 Entropy profile
Here the coefficient of the logarithm is due to the two disconnected Fermi seas $[q_{-},q_{+}]$ and $[-q_{+},-q_{-}]$ in terms of the reduced momentum, which is known to give a doubling for simple Toeplitz matrices [51, 52]. Note that, even though $q_{-}(\zeta)\to 0$ for $\zeta\to 0$ such that the two Fermi seas merge, the jump singularities still persist due to sign change in the offdiagonal elements in (12). As shown in Appendix B, the leading logarithmic term in (33) can be verified using recent results on block-Toeplitz determinants with Fisher-Hartwig singularities [53, 54, 55]. However, the non-universal piece $S_{0}(\zeta)$ in the entropy profile is given by the halved subleading constant, for which one has no explicit result available, and has to be extracted numerically.
@@ 5 Three-band model
Finally, let us have a look at the entropy profile, shown in Fig. 8. It has a a jump in the interior of the profile, which can be understood from the structure of the dispersion in Fig 6. Indeed, the three bands are now not identical, and the middle one delivers the global maximum $v_{1}$ of the velocity, whereas the bottom/top bands deliver a local maximum $v_{2}<v_{1}$. Hence, the outermost part $v_{2}<|\nu|<v_{1}$ of the profile corresponds to contributions from the middle band only. Indeed, by comparing the profiles at different times, one can verify that it is described by
@@
Figure 8: Entropy profiles at time $t=300$ in a trimerized chain with $N=200$ cells, as a function of $\nu=n_{0}/t$ and for different values of $\delta$. The dashed lines indicate the maxima of the quasiparticle velocities in the middle (red) and the upper/lower bands (blue).
@@ Appendix A Stationary phase calculation
The hydrodynamic limit corresponds to $m,n,t\gg 1$, with $m/t$ and $n/t$ kept fixed such that $|m-n|/t\ll 1$. In turn, this amounts to a stationary phase analysis of the double integral in (44), where the dominant contribution is due to the pole at $p=q$. Introducing the variables $Q=q-p$ and $P=(q+p)/2$, one can thus expand the oscillatory phases around $Q=0$ via the group velocity $v_{P}=-\omega^{\prime}_{P}$ as
@@
where $x=m+n-1$. This is precisely the result reported in (12)
@@ Appendix B Block Toeplitz determinants
The symbol of the block-Toeplitz matrix associated to the correlation matrix (12) has jumps at ±$\pm k_{\pm}$ (with $k_{\pm}=2q_{\pm}$ being the sublattice momentum), and the relevant limits of the symbol are obtained as
@@ Appendix C Particle number fluctuations for $\delta\to 1$
[…]
@@
Since this expression decays as $1/r^{2}$, after multiplying with $L-r$ the two pieces of the sum in (67) delivers a linear as well as a logarithmic contribution in $L$. The extensive term can be evaluated by exchanging $r=2j$, extending the sum to infinity and using
@@
Hence, taking the factor $-2$ in (65) into account, the extensive part exactly cancels with the diagonal sum in (66). Note, however, that if we are interested in the $\mathcal{O}(1)$ contributions to the fluctuations, one should keep also the $1/L$ corrections, which originate from extending the sum to infinity. Indeed, one has

### v2

@@ 2 Model and setup
[…]
@@ 3 Correlations and hydrodynamics
The result (12) has a very clear interpretation in terms of the quasiparticle picture. Indeed, the two terms simply correspond to contributions from the lower and upper band of quasiparticles $\alpha_{q}$ and $\beta_{q}$, respectively. Due to the symmetry of the dispersion in (4), these modes propagate ballistically with velocities $\pm v_{q}$, and contribute to the correlations if they reach the midpoint between the cells $m$ and $n$. Note also that (12) can be rewritten more compactly as
@@ 3.1 Density and current
We first evaluate the particle density and current defined in (8). Note that the result is going to depend only on $|\delta|$, and we assume $0<\delta<1$ to simplify notation. The maximal quasiparticle velocity can be obtained via the solution of $\frac{\mathrm{d}v_{q}}{\mathrm{d}q}=0$, which yields
@@ 4.1 CFT results
Having identified the inhomogeneous CFT, the next step is to apply the replica trick and twist-field machinery [48, 49] for the computation of the entropy of the domain $A=[-\infty,x]$, which has been carried out in [22]. The key is to map the inhomogeneous CFT (28), defined on a strip $0,\pi]\times\mathbb{R}$ in terms of the isothermal coordinates (27), into a flat one living on the upper half plane. This can be achieved by the subsequent application of a Weyl-transformation $\mathrm{e}^{\sigma(z,\bar{z})}$, and the exponential map $g(z)=\mathrm{e}^{iz}$. The twist field inserted at $(x,y)$ in the original geometry is thus mapped into $g(z)$ on the flat upper half plane, which amounts to calculating the entropy of a segment of length $\mathrm{Im}\,g(z)$ located at the boundary of a half-infinite chain. Taking into account the Jacobians of the conformal transformations, one obtains an effective length scale and the universal contribution to the entropy is given by [22]
@@
However, this is not the end of the story, as we have to deal with a lattice problem, which introduces a space-dependent non-universal contribution to the entropy. Nevertheless, this can be found by applying a local density approximation and using the known results for the constant term in a homogeneous chain [50]. Indeed, the time-evolved state is locally described by a boosted Fermi sea $k_{-},k_{+}]$ with boundaries $k_{\pm}=\frac{\pi}{2}\pm\arccos\nu$, such that the lattice contribution is given by
@@ 4.2 Entropy profile
Here the coefficient of the logarithm is due to the two disconnected Fermi seas $q_{-},q_{+}]$ and $q_{+},-q_{-}]$ in terms of the reduced momentum, which is known to give a doubling for simple Toeplitz matrices [51, 52]. Note that, even though $q_{-}(\zeta)\to 0$ for $\zeta\to 0$ such that the two Fermi seas merge, the jump singularities still persist due to sign change in the offdiagonal elements in (12). As shown in Appendix B, the leading logarithmic term in (33) can be verified using recent results on block-Toeplitz determinants with Fisher-Hartwig singularities [53, 54, 55]. However, the non-universal piece $S_{0}(\zeta)$ in the entropy profile is given by the halved subleading constant, for which one has no explicit result available, and has to be extracted numerically.
@@ 5 Three-band model
Finally, let us have a look at the entropy profile, shown in the left of Fig. 8, together with the corresponding density profile on the right. The entropy has a a jump in the interior of the profile, which can be understood from the structure of the dispersion in Fig 6. Indeed, the three bands are now not identical, and the middle one delivers the global maximum $v_{1}$ of the velocity, while the upper/lower bands deliver a local maximum $v_{2}<v_{1}$. Hence, the outermost part $v_{2}<|\nu|<v_{1}$ of the profile corresponds to contributions from the middle band only. The same feature can also be observed in the density, which develops a superimposed edge regime around $|\nu|=v_{2}$, where the fastest quasiparticles from the upper/lower bands start to contribute. Note also that the density develops a fine structure, with different hydrodynamical limits on the sublattices. This indicates that the analogue of eq. (12) must possess a more complicated structure, with non-uniform entries along the diagonals of the corresponding $3\times 3$ matrices in the integrals.
Comparing entropy profiles at different times, one can verify that it is described by
@@
Figure 8: Entropy (left) and density (right) profiles at time $t=300$ in a trimerized chain with $N=200$ cells, as a function of $\nu=n_{0}/t$ and for different values of $\delta$. The dashed lines indicate the maxima of the quasiparticle velocities $v_{1}$ in the middle (red) and $v_{2}$ in the upper/lower bands (blue).
@@ Appendix A Stationary phase calculation
The hydrodynamic limit corresponds to $m,n,t\gg 1$, with $m/t$ and $n/t$ kept fixed such that $|m-n|/t\ll 1$. In turn, this amounts to a stationary phase analysis of the double integral in (44), where the dominant contribution is due to the pole at $p=q$. Introducing the variables Q=qpQ=q-p and $P=(q+p)/2$, one can thus expand the oscillatory phases around $Q=0$ via the group velocity $v_{P}=-\omega^{\prime}_{P}$ as
@@
where x=m+n1x=m+n-1. This is precisely the result reported in (12)
@@ Appendix B Block Toeplitz determinants
The symbol of the block-Toeplitz matrix associated to the correlation matrix (12) has jumps at $\pm k_{\pm}$ (with $k_{\pm}=2q_{\pm}$ being the sublattice momentum), and the relevant limits of the symbol are obtained as
@@ Appendix C Particle number fluctuations for $\delta\to 1$
$\sum_{m<n\in A}|C_{m,n}|^{2}=$ $\sum_{\textrm{$r$ even}}^{L-2}(L-r)\mathcal{S}^{2}(r)+\sum_{\textrm{$r$ odd}}^{L-1}\left[\frac{L-r+1}{2}\mathcal{C}^{2}(r-1)+\frac{L-r-1}{2}\mathcal{C}^{2}(r+1)\right]$
$=$ $\sum_{\textrm{$r$ even}}^{L-2}(L-r)[\mathcal{S}^{2}(r)+\mathcal{C}^{2}(r)],$ (67)
@@
Since this expression decays as $1/r^{2}$, after multiplying with LrL-r the two pieces of the sum in (67) delivers a linear as well as a logarithmic contribution in $L$. The extensive term can be evaluated by exchanging $r=2j$, extending the sum to infinity and using
@@
Hence, taking the factor 2-2 in (65) into account, the extensive part exactly cancels with the diagonal sum in (66). Note, however, that if we are interested in the $\mathcal{O}(1)$ contributions to the fluctuations, one should keep also the $1/L$ corrections, which originate from extending the sum to infinity. Indeed, one has